### v1

# Stochastic synaptic dynamics under learning

Abstract
Learning is based on synaptic plasticity, which drives and is driven by neural activity. Because pre- and postsynaptic spiking activity is shaped by randomness, the synaptic weights follow a stochastic process, requiring a probabilistic framework to capture the noisy synaptic dynamics. We consider a paradigmatic supervised learning example: a presynaptic neural population impinging in a sequence of episodes on a recurrent network of integrate-and-fire neurons through synapses undergoing spike-timing-dependent plasticity (STDP) with additive potentiation and multiplicative depression. We first analytically compute the drift- and diffusion coefficients for a single synapse within a single episode (microscopic dynamics), mapping the true jump process to a Langevin and the associated Fokker-Planck equations. Leveraging new analytical tools, we include spike-time–resolving cross-correlations between pre- and postsynaptic spikes, which corrects substantial deviations seen in standard theories purely based on firing rates. We then apply this microdynamical description to the network setup in which hetero-associations are trained over one-shot episodes into a feed-forward matrix of STDP synapses connecting to neurons of the recurrent network (macroscopic dynamics). By mapping statistically distinct synaptic populations to instances of the single-synapse process above, we self-consistently determine the joint neural and synaptic dynamics and, ultimately, the time course of memory degradation and the memory capacity. We demonstrate that specifically in the relevant case of sparse coding, our theory can quantitatively capture memory capacities which are strongly overestimated if spike-time–resolving cross-correlations are ignored. We conclude with a discussion of the many directions in which our framework can be extended.
I Introduction
Synaptic plasticity constitutes the foundation of learning. Although the mechanisms responsible for plasticity are diverse and incompletely understood [1, 2], the quantitative dependence of plasticity on the relative timing between pairs of pre- and postsynaptic spikes [3, 4, 5, 6] accounts for many of the experimental observations; corrections are often extensions of this pair-based rule. The statistics of synaptic weight dynamics is thus determined by the statistics of pre- and postsynaptic spikes. Because the latter display a considerable variability, synaptic dynamics has to be described in a stochastic framework, even if the intrinsic noise sources of plasticity [7, 8, 9, 10] are neglected.
Several aspects of synaptic dynamics have been studied before; the mean dynamics and even ensemble dynamics [11, 12, 13, 14, 15, 16, 17, 18, 19] have been investigated for a variety of plasticity rules and settings, see [2] for an extensive review. More recently, the conjunction of classical spike-timing-dependent plasticity (STDP) with homeostatic plasticity has received increased attention (see [20] for an overview).
Classical pair-based STDP implies that the expected synaptic weight depends on (i) pre- and postsynaptic firing rates and (ii) cross-correlations between pre- and postsynaptic spikes [11]. Since the weight change induced by STDP is sensitive to relative spike-timing differences on the order of milliseconds [4, 5, 1], these two contributions play fundamentally distinct roles for learning. The effect of cross-correlations on learning is typically taken into account by interpreting the post-synaptic neuron as a linear filter. This somewhat drastic simplification is either explained by model choice (conditionally Poissonian neuron) or through a daring but educated trick: to linearize nonlinear neuron models per realization using realization-averaged response functions [21, 22, 23], see e.g. [24, 25, 17]. Response functions can be derived with Fokker-Planck theory, for integrate-and-fire neurons see Refs. [26, 27, 28, 29, 30, 31].
Despite the vast body of research on synaptic dynamics, a robust and generalizable stochastic description of synaptic dynamics is still missing. Such a description is clearly needed to develop multiscale memory models that take into account the single neuron level and the population level. In this paper we lay the foundations to this endeavor; specifically, we analytically characterize the stochastic process of synapses endowed with STDP that feed Poisson spikes to leaky integrate-and-fire (LIF) neurons and are shaped by the spikes of both Poisson and LIF neurons. We focus on the stochasticity due to pre- and postsynaptic fluctuations and neglect the intrinsic noisiness of synapses. Instead of relying on realization-wise linearization, we incorporate cross-correlations through an exact model-independent relation between cross-correlations and response-functions that two of us reported recently [32]. This approach is robust and generalizable.
Synapses endowed with STDP are capable of self-organized formation of diverse types of memory including (i) learning input correlations [33], (ii) reinforcement learning in recurrent networks of LIF neurons [17] or (iii) error correction of drifting assemblies [34]. Yet another type of STDP-based memory are hetero-associations, i.e., representations in one neuron population evoke representations in another one. Building on our results about stochastic synaptic dynamics, we analytically investigate hetero-associative memory. As we outline, for characterizing this type of memory, the stochastic view on synaptic dynamics is indispensable. Extending our methods to the aforementioned mechanisms (i–iii) poses exciting problems for future work.
The paper is organized as follows: In Sec. II, we first introduce the synapse and neuron model Fig. 1(a,b) studied throughout the paper. Building on these models, we introduce the network model Fig. 1(c) and the training scheme Fig. 1(d), which store hetero-associations into a plastic feed-forward matrix. In Sec. III, we characterize the stochastic process of a single synaptic weight by deriving its drift- and diffusion coefficients, which define a Langevin (or corresponding Fokker-Planck) equation. From this description, we derive the dynamics of the ensemble mean and variance of synaptic weights. In Sec. IV, we leverage the theory of single weights to study the network scenario in a mean-field theory of synaptic populations. We apply the mean-field theory to characterize forgetting and compute the memory capacity of the setup.
Figure 1: We investigate memory properties by studying a single synapse dynamics that captures network properties effectively by a mean-field theory. (a) A Poisson process $\eta$ drives the neuron $x$ through the synapse $w$. The weight $w$ is plastic, i.e., it dynamically depends on $x$ and $\eta$. (b) STDP window Eq. (2) for $w=0.01$ (light gray) through $w=0.2$ (black). The setup (a,b) is studied in Sec. III (c) A cue representation $\boldsymbol{q}$ in a layer of Poisson processes (gray) is activated and drives a recurrent neural network through a matrix of plastic synapses. Among the excitatory neurons (red) in the recurrent network, a target representation $\boldsymbol{p}$ receives an additional stimulus. (d) Training procedure to store hetero-associations between cues and targets. The setup (c,d) is studied in Sec. IV. Parameters $\Delta_{c}=2\times 10^{-3}$, $r_{ac}=8\times 10^{-3}$, $\tau_{c}=16.8/20$, and $\tau_{ac}=33.7/20$ are matched to [5].
II Models
First, we introduce the synaptic model depicted in Fig. 1(a,b). This model depends on pre- and postsynaptic spike trains, thus, next, we define the neuron models used in this paper. Then we introduce the network model depicted in Fig. 1(c), which is composed of the synapse- and neuron models above. We are interested how hetero-associations can be stored into this network; the necessary training scheme [see Fig. 1(d)] is explained last.
II.1 Synapse model
We consider a synapse with presynaptic spike train $\eta(t)=\sum_{i}\delta(t-t_{i}^{\text{pre}})$ and postsynaptic spike train $x(t)=\sum_{j}\delta(t-t_{j}^{\text{post}})$. The evolution of the synaptic weight $w$ in the classical STDP model [2] can be described by
$\dot{w}(t)$ $=\int_{-\infty}^{t}dt^{\prime}\,\kappa[t-t^{\prime},w(t)]\eta(t^{\prime})x(t)$
$\;+\int_{-\infty}^{t}dt^{\prime}\,\kappa[t^{\prime}-t,w(t)]\eta(t)x(t^{\prime}).$ (1)
Each pre-post spike pair leads to a jump in $w$ by $\kappa$; causal spike pairs are captured by the first line, anti-causal pairs by the second line. Here, $\kappa(\tau,w)$ is the STDP window function, which we assume only depends on the time difference between spikes and on $w(t)$. The lower integral bounds are set to -$\infty$ so as to avoid switch-on effects of the plasticity rule. Throughout the paper, we consider an exponential window with multiplicative (i.e., $w$-dependent) depression and additive (i.e., $w$-independent) potentiation [12, 35]
$\kappa(\tau,w)=\begin{cases}\Delta_{c}e^{-\tau/\tau_{c}}&\tau\geq 0\\ -r_{ac}we^{\tau/\tau_{ac}}&\tau<0,\end{cases}$ (2)
see Fig. 1(b), with $\Delta_{c},r_{ac}>0$ determining the amplitudes and $\tau_{c},\tau_{ac}>0$ being the time scales of potentiation and depression, respectively. In addition to Eq. (1) and Eq. (2), we demand that the considered excitatory synapse maintains a positive weight throughout, $w(t)\geq 0$, a necessary condition for Dale’s law, and implemented in our simulations by a clipping boundary condition. The plasticity rule Eq. (2) with its multiplicative depression and additive potentiation is chosen as it captures the experimental findings of Ref. [5] and because of its inherent stability. Generalizations of Eq. (1), e.g., triplet rules [36, 37] or neuromodulatory dynamics [19] are of interest but out of scope of our study.
For the exponential window Eq. (2), one can rewrite Eq. (1) in terms of pre- and postsynaptic trace variables $A_{\text{pre}}$ and $A_{\text{post}}$ [38]
$\dot{w}$ $=\Delta_{c}A_{\text{pre}}x-r_{ac}wA_{\text{post}}\eta,$
$\dot{A}_{\text{pre}}$ $=-\tau_{c}^{-1}A_{\text{pre}}+\eta,$
$\dot{A}_{\text{post}}$ $=-\tau_{ac}^{-1}A_{\text{post}}+x.$ (3)
This embedding is useful for simulations as it is makes the system local in time. To avoid switch-on effects in Eq. (3), we integrate $A_{\text{pre}}$ and $A_{\text{post}}$ for some time before we start integrating $w$; this way, $A_{\text{pre}}$ and $A_{\text{post}}$ thermalize, i.e., their statistics become stationary, which is in line with the lower integral bound at -$\infty$ in Eq. (1).
II.2 Neuron model
We here introduce a neuron model that is driven (i) by Gaussian white noise and (ii) by a single presynaptic neuron through a single synapse. The dynamics of this synapse will be studied in Sec. III. The dynamics of multiple synapses providing input to a neuron, studied in Sec. IV, can be effectively mapped to the single-synapse case introduced here. The synaptic dynamics Eq. (1) is driven by the pre- and postsynaptic neural activity $\eta$ and $x$, respectively. Thus, to specify the synaptic dynamics, we need to specify the neuron models. Throughout the paper, $\eta(t)$ is the spike train of a Poisson process with rate $\nu$, and $x(t)$ is the spike train of a model neuron driven by $\eta$. The developed framework holds true for arbitrary neuron models $x$, but we give specific expressions and simulation results for the LIF neuron $x(t)=\sum_{i}\delta(t-t_{i}^{\text{post}})$, where $t_{i}^{\text{post}}$ are the times at which the membrane voltage governed by
$\tau_{m}\dot{v}=-v+\mu+\sqrt{2D}\xi(t)+w(t)\eta(t)$ (4)
hits the threshold $v_{\text{t}}$; additionally, at $t_{i}^{\text{post}}$, $v$ is reset to $v_{\text{r}}$. Throughout the manuscript, $v_{\text{r}}=0$, $v_{\text{t}}=1$, and the membrane time constant is set to $\tau_{m}=1$, i.e., we measure time in multiples of $\tau_{m}$. The input to Eq. (4) consists of the mean current $\mu$, white Gaussian noise $\sqrt{2D}\xi$ with noise intensity $D$ and $\left\langle\xi(t)\xi(t^{\prime})\right\rangle=\delta(t-t^{\prime})$, and the Poissonian spikes weighted by the synapse $w$. The LIF neuron is mechanistic enough to generate spike trains $x$ that are correlated with $\eta$ in a beyond-rate-based fashion. Moreover, it has been shown to reproduce experimental spike trains of pyramidal cells [39], and many of its statistical properties are available [40, 26, 27, 41].
II.3 Network model
Here, we introduce the network model depicted in Fig. 1(c), which can store hetero-associations, i.e., feed-forward associations between representations in two distinct populations. A layer of $M$ Poisson processes $\eta_{i}$ (gray circles) drives a recurrent network of $N$ LIF neurons [42] (red and blue circles) through a plastic weight matrix $W$. The recurrent network consists of $N_{E}$ excitatory and $N_{I}$ inhibitory neurons, among which only the excitatory neurons are targeted by $W$. Addidionally, each neuron has exactly $C_{E}$ incoming excitatory connections from the recurrent network and $C_{I}$ incoming inhibitory connections with static weights $J_{EE}=J$, $J_{EI}=J_{II}=-gJ$, $J_{IE}=hJ$, parameterized by $J>0$, $g>0$, and $h>0$.
Most input Poisson processes have a low firing rate $\nu_{\text{lo}}$, however, in each training step, a subset of $f_{c}M$ Poisson generators, represented by the binary vector $\boldsymbol{q}\in\left\{0,1\right\}^{M}$ fires with increased rate $\nu_{\text{hi}}$. Simultaneously, a subset of $f_{s}N_{E}$ excitatory neurons, represented by the binary vector $\boldsymbol{p}\in\left\{0,1\right\}^{N_{E}}$, receives direct Poissonian input $\gamma_{i}$ with rate $\nu_{s}$ and static weight $J_{s}$; the other $(1-f_{s})N_{E}$ excitatory neurons do not receive additional input. We are particularly interested in sparse representations, where the activation ratios $f_{c},f_{s}\ll 1$.
Summarizing the above, the membrane voltage of excitatory neurons follows
$\dot{v}_{i}$ $=-v_{i}+\mu_{E}+\sqrt{2D_{E}}\xi_{i}(t)+J\sum_{n\in C_{E}(i)}x_{n}-gJ\sum_{n\in C_{I}(i)}x_{n}$
$\quad+\sum_{j}W_{ij}(t)\eta_{j}(t)+p_{i}J_{s}\gamma_{i}(t),$ (5)
and the membrane voltage of inhibitory neurons obeys
$\dot{v}_{i}=-v_{i}+\mu_{I}+\sqrt{2D_{I}}\xi_{i}(t)+hJ\sum_{n\in C_{E}(i)}x_{n}-gJ\sum_{n\in C_{I}(i)}x_{n},$ (6)
where the spike trains $x_{n}(t)=\sum_{i}\delta(t-t_{n,i})$ are given in terms of the fire-and-reset times $t_{n,i}$ of neuron $n$ [see Eq. (4)], and $\mu_{E},D_{E}$ ($\mu_{I},D_{I}$) are the baseline mean input and noise-intensity of the excitatory (inhibitory) neurons. The recurrent network is a variant of [42, Model A with $h=1$]. For $h\neq 1$, excitatory spikes have different efficacies at excitatory and inhibitory neurons, respectively. For $h>1$, this establishes a competition between the neurons that proves useful for long memory. In a study of rat neocortex, $J_{IE}\approx 2J_{EE}$ (i.e., $h=2$) has been reported for regular spiking excitatory neurons and fast spiking inhibitory interneurons [43, 44].
II.4 Training scheme
The training procedure described here mimics the situtation where a pre-synaptic assembly (a cue) is activated and a supervisor drives a target assembly in the post-synaptic population. Then, through STDP, the matrix $W$ learns to associate the cue with the target: after some time, activating the cue will autonomously activate the target without requiring the supervisor—the hetero-association is stored.
Specifically, the goal of training is to store associations of random (independently drawn) pattern pairs $(\boldsymbol{q}_{k},\boldsymbol{p}_{k})$ into the network Fig. 1(c). The training procedure is illustrated in Fig. 1(d) and proceeds as follows. The zeroth association is trained by setting $\boldsymbol{q}=\boldsymbol{q}_{0}$ and $\boldsymbol{p}=\boldsymbol{p}_{0}$. First, trace variables and membrane voltages are integrated for $T_{\text{warm}}=20$ with frozen $W$. Second, all dynamical variables, including the entrees of $W$, are integrated for a time $T$; here, the synapses $W_{ij}$ follow the dynamics Eq. (1) with the specific kernel Eq. (2), see Appendix A for implementation details. We assume that there is a pause before the next pattern pair is trained. During this pause, homeostatic plasticity occurs, which, according to experimental findings, can be modeled as a slow rescaling of weights to maintain firing rates (see [45, 20] and references therein) or summed synaptic weight per postsynaptic neuron (approximating the experimentally observed conservation of summed synapse surface area per postsynaptic neuron over time, see [46]). We follow the latter view, and thus, after each training session, we rescale each weight as
$W_{ij}\rightarrow W_{ij}\frac{m_{0}}{M^{-1}\sum_{j^{\prime}=1}^{M}W_{ij^{\prime}}},$ (7)
where we introduced the parameter $m_{0}$ that defines the average synaptic weight per postsynaptic neuron. Note that each row of $W$ sums to $Mm_{0}$. After homeostasis, we set $\boldsymbol{q}=\boldsymbol{q}_{1}$ and $\boldsymbol{p}=\boldsymbol{p}_{1}$, and repeat the entire procedure, then proceed with $(\boldsymbol{q}_{2},\boldsymbol{p}_{2})$ and so on.
Figure 2: Stochastic dynamics of single synapses. (a1) Sample trajectories of Eq. (1) and (a2) sample trajectories of the corresponding Langevin equation for $\mu=0.6$, $D=0.2$, $\nu=0.1$, $m_{0}=0.1$, $\sqrt{V_{0}}=10^{-3}$ and STDP parameters as in Fig. 1(b). (b1–b4) Drift coefficient $D^{(1)}$ from simulations (orange circles) and theory [Eq. (17) (black line) and single contributions: firing-rate (gray), mean-response (green), and noise-intensity–response (blue)] for different $\mu$ and $D$. The instantaneous firing rate $r$ at $w=0.1$ is indicated in the upper center of the four panels. (c1–c4) Finite-time diffusion coefficient $\Delta t=10$. Simulation results (orange circles). Theory (black lines) $V(\Delta t)+(m(\Delta t)-w)^{2}]/(2\Delta t)$, with $m$ from Eq. (24) and $V$ from Eq. (26).
III Stochastic synaptic dynamics
The dynamics of a synaptic weight $w$, Eq. (1), is stochastic due to the randomness of $\eta$ and $x$. Trajectories of $w$ can be obtained in simulations, for instance, for the window function Eq. (2), one has to integrate both stochastic equations, Eq. (3) and Eq. (4). The synaptic weight performs a jump process Fig. 2(a1): with each new pre- or post-synaptic spike, a new set of spike pairs is formed that leads to a finite jump, and between jumps the weight is constant. In this section, we develop a simplified description in terms of a Langevin equation and use it to calculate the transient mean and the variance of an ensemble of synapses.
The rate of change of synaptic weight is given by Eq. (1). Assuming that the amplitudes $\kappa$ are sufficiently small and the rate of spike pairs sufficiently high, $\dot{w}$ can be approximated by a white Gaussian process. This leads to a Langevin equation
$\dot{w}=D^{(1)}(w)+\sqrt{2D^{(2)}(w)}\zeta(t),$ (8)
which, due to causality of the synaptic updates, we interpret in the sense of Ito [47]. Here, $\zeta(t)$ is white Gaussian noise with $\left\langle\zeta(t)\right\rangle=0$ and $\left\langle\zeta(t)\zeta(t^{\prime})\right\rangle=\delta(t-t^{\prime})$. Samples of Eq. (8) are shown in Fig. 2(a2). Correspondingly, the transition probability $p(w,t|w_{0},t_{0})$ of weights follows the Fokker-Planck equation
$\partial_{t}p(w,t|w_{0},t_{0})$ $=-\partial_{w}D^{(1)}(w)p(w,t|w_{0},t_{0})$
$\quad+\partial_{w}^{2}D^{(2)}(w)p(w,t|w_{0},t_{0}).$ (9)
In this section, we derive the functions $D^{(1)}(w)$ and $D^{(2)}(w)$, which we then apply to describe the evolution of an ensemble of synaptic weights.
III.1 Drift coefficient
The drift coefficient $D^{(1)}$ is the first Kramers-Moyal coefficient
$D^{(1)}(w)=\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\left\langle w_{\text{traj}}(t+\Delta t)-w_{\text{traj}}(t)\right\rangle_{w_{\text{traj}}(t)=w},$ (10)
where $w_{\text{traj}}$ is a sample of the stochastic process Eq. (1), and the subscript denotes a condition on the ensemble average. Assuming slow weight dynamics (compared to the neuron’s thermalization time scale), Eq. (10) can be generally evaluated
$D^{(1)}(w)=\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)\left[\nu r(w)+C_{x\eta}(\tau,w)\right],$ (11)
which is a well known result [11]. Here, $r(w)\equiv\left\langle x\right\rangle$ is the instantaneous firing rate of the postsynaptic neuron assuming weight $w$, and $C_{x\eta}(\tau,w)=\left\langle x(t+\tau)\eta(t)\right\rangle_{w}-r(w)\nu$ is the input-spikes–output-spikes cross-correlation. Within a diffusion approximation of the input to the LIF neuron
$\mu+\sqrt{2D}\xi(t)+w\eta(t)$
$\approx$ $\mu_{\text{DA}}(\nu)+\sqrt{2D_{\text{DA}}(\nu)}\xi(t),$ (12)
where $\mu_{\text{DA}}(\nu)=\mu+w\nu$ and $D_{\text{DA}}(\nu)=D+w^{2}\nu/2$, the instantaneous firing rate is given by [40]
$r^{-1}=\int_{\frac{v_{r}-\mu_{\text{DA}}}{\sqrt{2D_{\text{DA}}}}}^{\frac{v_{t}-\mu_{\text{DA}}}{\sqrt{2D_{\text{DA}}}}}ds\,e^{s^{2}}[1+\text{erf}(s)],$ (13)
see [48] for an efficient and stable evaluation of Eq. (13). For the STDP kernel in Eq. (2), the first, firing-rate–based, part in Eq. (11) is
$\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)\nu r(w)=(\Delta_{c}\tau_{c}-r_{ac}w\tau_{ac})\nu r(w).$ (14)
The interesting part in Eq. (11) from a spike-coding perspective is the cross-correlation $C_{x\eta}$. Since $\eta$ is assumed to be Poissonian, the cross-correlation is (despite the nonlinearity of the neuron model) exactly related to the firing-rate response [32]
$C_{x\eta}(\tau)=\nu\frac{\delta}{\delta\nu(t)}\left\langle x(t+\tau)\right\rangle,$ (15)
where $\frac{\delta}{\delta\nu(t)}f[\nu]\equiv\lim_{h\rightarrow 0}\frac{d}{dh}f[\nu+h\delta(t-\circ)]$ denotes a functional derivative (here, $\circ$ represents the time argument of $\nu$ in $f$). The response function of the output spikes of a LIF neuron to rate modulations is to our knowledge not known (but see [29] for such a result if the input amplitudes are exponentially distributed). However, in the diffusion approximation of the LIF neuron’s input in Eq. (12), one can apply the chain rule to obtain
$\frac{\delta\left\langle x(t+\tau)\right\rangle}{\delta\nu(t)}\approx w\frac{\delta\left\langle x(t+\tau)\right\rangle}{\delta\mu_{\text{DA}}(t)}+\frac{1}{2}w^{2}\frac{\delta\left\langle x(t+\tau)\right\rangle}{\delta D_{\text{DA}}(t)}.$ (16)
Thus, the response function in Eq. (15) can be approximated in terms of the response functions to mean- and to noise-intensity modulations. Their Fourier transforms, $\alpha(\Omega)$ (the susceptibility to mean modulations) and $\beta(\Omega)$ (the susceptibility to noise-intensity modulations), can be derived with Fokker-Planck theory and are known in terms of special functions [26, 27], also presented in Appendix B. For the exponential STDP kernel in Eq. (2), the integral in Eq. (11) depends on the Laplace transform of the response functions evaluated at $1/\tau_{c}$ [note that due to causality, $C_{x\eta}(\tau<0)=0$]; since the Fourier transforms $\alpha$ and $\beta$ are analytic, we may simply evaluate them at $i/\tau_{c}$. Thus, summing up, the drift coefficient is
$D^{(1)}(w)$ $=(\Delta_{c}\tau_{c}-r_{ac}w\tau_{ac})\nu r$
$\;+\Delta_{c}\nu\left[w\alpha(i\tau_{c}^{-1})+\frac{1}{2}w^{2}\beta(i\tau_{c}^{-1})\right].$ (17)
$D^{(1)}$ is shown and dissected into different contributions in Fig. 2(b) and shows agreement with simulation results of the system. For small weights, the drift is dominated by the firing-rate contributions (gray lines), whereas for increasingly larger weights the response contributions dominate the drift. The mean-response contribution $\Delta_{c}\nu w\alpha(i\tau_{c}^{-1})$ and the noise-intensity–response contribution $\Delta_{c}\nu(1/2)w^{2}\beta(i\tau_{c}^{-1})$ are shown in Fig. 2(b) by the green and the blue line, respectively. When computing the drift using realization-wise linearization, the noise-intensity response is missing. We have not encountered severe violations of Eq. (17) in any reasonable regime. For other neuron models than the LIF neuron, one only needs to substitute the respective mean- and noise-intensity responses.
III.2 Diffusion coefficient
The diffusion coefficient is the second Kramers-Moyal coefficient
$D^{(2)}(w)$ $=\frac{1}{2}\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}$
$\times\left\langle\left[w_{\text{traj}}(t+\Delta t)-w_{\text{traj}}(t)\right]^{2}\right\rangle_{w_{\text{traj}}(t)=w},$ (18)
where the average is over sample trajectories $w_{\text{traj}}$ of Eq. (1). Since the weight involves integrals over the product of $x$ and $\eta$, the squared weights in Eq. (18) lead to four-point correlation functions. This makes it difficult to exactly evaluate Eq. (18) in general. However, the diffusion on time scales $\Delta t$ above the inverse firing rates may be estimated by applying Wick’s theorem, as an approximation since $(\eta,x)$ is a on-Gaussian process, to the four-point correlation functions, which leads to (see Appendix C for details)
$D^{(2)}(w)\approx\frac{1}{4}r\nu(\Delta_{c}^{2}\tau_{c}+r_{ac}^{2}w^{2}\tau_{ac}).$ (19)
Eq. (19) is the infinitesimal diffusion coefficient. For non-infinitesimal $\Delta t$ long enough (specifically, in Fig. 2(c), $\Delta t=10$), the diffusion is captured by the Langevin dynamics based on Eq. (17) and Eq. (19), as we derive in the next section.
III.3 Mid- and long-term evolution of an ensemble of synapses
The drift- and diffusion coefficients derived above describe the process on infinitesimal times. Here, we study the mean and variance of an ensemble of synapses on finite times. We provide a technically detailed analysis because the results of this section are the main building blocks of the mean-field theory of learning discussed in Sec. IV. Although the dynamics can admit a stationary solution [see e.g. the zero of $D^{(1)}(w)$, i.e., the black solid line crosses zero in Fig. 2(b3-b4)], we are here mainly interested in transient dynamics: If the training patterns permanently (but slowly) change and the synapse is subject to slow homeostatic plasticity, the stationary state will never be reached in a learning situation. Therefore we here study the non-equilibrium dynamics of the ensemble mean and variance.
We start the discussion from a general point of view and recover that the drift- and diffusion coefficients derived above suffice for our purpose. The transition probability $p$ of a Markov process obeys the Kramers-Moyal expansion
$\frac{\partial}{\partial t}p(w,t|w^{\prime},t^{\prime})=\sum_{n=1}^{\infty}\left(-\frac{\partial}{\partial w}\right)^{n}\left[D^{(n)}(w)p(w,t|w^{\prime},t^{\prime})\right],$ (20)
with the Kramers-Moyal coefficients
$D^{(n)}(w)$ $=\frac{1}{n!}\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}$
$\times\left\langle\left[w_{\text{traj}}(t+\Delta t)-w_{\text{traj}}(t)\right]^{n}\right\rangle_{w_{\text{traj}}(t)=w}.$ (21)
When assuming vanishing probability density at the boundaries, a consequence of Eq. (20) is [49]
$\partial_{t}\left\langle w^{n}\right\rangle=\sum_{k=1}^{n}\frac{n!}{(n-k)!}\left\langle w^{n-k}D^{(k)}(w)\right\rangle.$ (22)
The Langevin approximation, which only considers the first two Kramers-Moyal coefficients, thus implies
$\dot{m}$ $=\left\langle D^{(1)}(w)\right\rangle$
$\dot{V}$ $=2\left[\left\langle wD^{(1)}(w)\right\rangle-\left\langle w\right\rangle\left\langle D^{(1)}(w)\right\rangle\right]+2\left\langle D^{(2)}(w)\right\rangle,$ (23)
where $m=\left\langle w\right\rangle$ and $V=\left\langle(w-\left\langle w\right\rangle)^{2}\right\rangle$ are the ensemble mean and variance. One may Taylor expand $D^{(1)}$ and $D^{(2)}$ around the self-consistent solution $m(t)$; if $D^{(1)}$ is sufficiently smooth, this leads to
$\dot{m}$ $=D^{(1)}(m)$ (24)
$\dot{V}$ $=2D^{(1)\prime}(m)V+\frac{r(m)\nu}{2}\left[\Delta_{c}^{2}\tau_{c}+r_{ac}^{2}\tau_{ac}\left(V+m^{2}\right)\right].$ (25)
We evaluate the mean dynamics Eq. (24) with a Runge-Kutta scheme; Eq. (25) can then be integrated and yields
$V$ $=e^{A(t)}\left(V_{0}+\int_{0}^{t}dt^{\prime}2D^{(2)}[m(t^{\prime})]e^{-A(t^{\prime})}\right)$
$A(t)$ $=\int_{0}^{t}2D^{(1)\prime}[m(t^{\prime})]dt^{\prime}+\frac{t}{2}r\nu r_{ac}^{2}\tau_{ac}.$ (26)
Eq. (24) and Eq. (26) can be efficiently evaluated to obtain the mean and variance of an ensemble of synapses. These equations are thus useful formulations of the microscopic (i.e., single-synapse) model, which we leverage in the next section to derive macroscopic (i.e., population) dynamics. We also use Eq. (24) and Eq. (26) to verify the second moment of the stochastic increment over a finite time window Fig. 2(c); i.e., due to the invalidity of the infinitesimal diffusion coefficient, we predict the finite-time diffusion coefficient as put forward in [50].
We lastly add a remark on the truncation of the stochastic process to only two Kramers-Moyal coefficients (Fokker-Planck level). To this end, we reflect on Eq. (22). While explicitly the evolution of the $n$th moment only depends on the first $n$ Kramers-Moyal coefficients, it implicitly depends on the full expansion through the probability density defining the expectation value. However, when the Kramers-Moyal coefficients are smooth enough and the spread of the ensemble small enough such that $D^{(n)}$ is well captured by its $n$th order Taylor approximation, then the evolution of the $n$th moment only depends on moments up to order $n$. Consequently, in this case we achieve moment closure at order $n$, i.e. the system is self-consistently described by the dynamics of the first $n$ moments.
IV Dynamics of learning
In this section we study the network and learning scheme sketched in Fig. 1(c,d) and introduced in Sec. II.3 and Sec. II.4 in Sec. II. The dynamics can now be viewed on two time scales: First, on the discrete macroscopic time scale (pattern index $k$, going from one training session to the next), the data that the network is exposed to is exchanged. Since subsequent pattern pairs are uncorrelated, the sequence of matrices $W_{k}$ after the $k$th training session including homeostasis is a Markov chain (see also [51]). Second, on the continuous microscopic time scale (within a training session), the present pattern pair $(\boldsymbol{q}_{k},\boldsymbol{p}_{k})$ appears static, and the fast neural and synaptic dynamics evolve as defined in Sec. II. As we discuss in detail in this section, the system becomes stationary on the macroscopic time scale while remaining non-equilibrium on the microscopic time scale. The following analysis starts by studying the storage of a memory, namely the association $\boldsymbol{q}_{k}\rightarrow\boldsymbol{p}_{k}$. This storage occurs on the microscopic time scale by exposing the system to $(\boldsymbol{q}_{k},\boldsymbol{p}_{k})$. The analysis proceeds by quantifying the stationary regime on the macroscopic time scale. Assuming this regime, we lastly study how memories are forgotten on the macroscopic time scale.
IV.1 Synaptic dynamics on the microscopic time
Here, we characterize the change of the synaptic weight matrix $W(t)$ during training session $k$. The starting point is the homeostatically scaled matrix from the previous session $W(0)=W_{k-1}$, where, as mentioned in the discussion of Eq. (7), each row of $W(0)$ sums to $Mm_{0}$. The new pattern pair $(\boldsymbol{q}_{k},\boldsymbol{p}_{k})$ defines four synaptic populations,
$P_{ab}=\{W_{ij}:\,p_{k,i}=a,\,q_{k,j}=b\}.$ (27)
These sets sort synapses by where they point to ($a=0$: to non-target neurons, $a=1$: to target neurons) and where they come from ($b=0$: from non-cue neurons, $b=1$: from cue neurons). Since $W(0)$ is still uncorrelated to $(\boldsymbol{q}_{k},\boldsymbol{p}_{k})$, the weights in all populations can be seen as randomly sampled from the set of weights in $W_{k-1}$. We describe the synaptic dynamics through the population means and variances
$\hat{m}_{ab}=\frac{1}{|P_{ab}|}\sum_{w\in P_{ab}}w,\quad\hat{V}_{ab}=\frac{1}{|P_{ab}|}\sum_{w\in P_{ab}}(w-\hat{m}_{ab})^{2},$ (28)
where $|\cdot|$ here denotes the number of elements in the sets Eq. (27). Strictly speaking, $\hat{m}_{ab}$ and $\hat{V}_{ab}$ are stochastic processes that depend on the realizations of the Poissonian and Gaussian processes fed into the network. Neglecting correlations, according to the central limit theorem one would estimate their fluctuations to scale as $1/\sqrt{|P_{ab}|}$, thus, if populations are large enough, it is reasonable to assume self-averaging $\hat{m}_{ab}\approx\left\langle\hat{m}_{ab}\right\rangle\equiv m_{ab}$ and $\hat{V}_{ab}\approx\left\langle\hat{V}_{ab}\right\rangle\equiv V_{ab}$, which strongly simplifies the analysis. The initial means are $m_{ab}(0)=m_{0}$ due to homeostasis, the initial variances $V_{ab}(0)=V_{0}$, where here $V_{0}$ should be understood as a parameter; later we fix it self-consistently to the stationary total variance of $W_{k\rightarrow\infty}$.
Our aim is to obtain $m_{ab}(t)$ and $V_{ab}(t)$ similarly to the single synapse case in Sec. III. To this end, we first map the neurons in the network Eq. (5) to instances of the single neuron Eq. (4), who’s input is a mean current and Gaussian white noise with fixed intensity. Thus, we need to calculate the mean and noise intensity of the input to the neurons. The input from the Poisson layer to the neurons in the target population ($a=1$) or in the non-target population ($a=0$) is determined by the first- and second-order statistics of Poisson processes
$\mu_{a}^{\text{cue}}$ $=M[m_{a1}f_{c}\nu_{\text{hi}}+m_{a0}(1-f_{c})\nu_{\text{lo}}]$
$D_{a}^{\text{cue}}$ $=\frac{1}{2}M[(m_{a1}^{2}+V_{a1})f_{c}\nu_{\text{hi}}+(m_{a0}^{2}+V_{a0})(1-f_{c})\nu_{\text{lo}}].$ (29)
The total input (including cue, recurrent network, and supervision) to neurons in population $a$ is thus by Poisson approximation as in [42]
$\mu_{a}^{\text{tot}}$ $=\mu_{E}+JC_{E}r_{E}-gJC_{I}r_{I}+\mu_{a}^{\text{cue}}+\delta_{a1}J_{s}\nu_{s},$
$D_{a}^{\text{tot}}$ $=D_{E}+\frac{1}{2}[J^{2}C_{E}r_{E}+(gJ)^{2}C_{I}r_{I}D_{a}^{\text{cue}}+\delta_{a1}J_{s}^{2}\nu_{s}],$ (30)
where $\delta_{a1}$ is the Kronecker symbol and $r_{E}$ and $r_{I}$ are the mean firing rates of the excitatory and the inhibitory neurons, respectively. These are determined by mean-field theory along the lines of [42], see Appendix D.
As one can see from the total effective input Eq. (30), the postsynaptic neurons are, within the employed approximation, effectively decoupled apart from their common dependence on the mean fields $m_{ab}$, $V_{ab}$, $r_{E}$, and $r_{I}$, which we now determine self-consistently. Thus, assuming knowledge of the numerical values of the mean fields, the single-neuron statistics is readily determined and we can proceed as in Sec. III: The drift- and diffusion coefficients of the four distinct synaptic populations are, respectively
$D_{ab}^{(1)}(w)$ $=(\Delta_{c}\tau_{c}-r_{ac}w\tau_{ac})\nu_{b}r_{a}(m,V)$
$\;+\Delta_{c}\nu_{b}\left[w\alpha_{a}(i\tau_{c}^{-1},m,V)+\frac{1}{2}w^{2}\beta_{a}(i\tau_{c}^{-1},m,V)\right],$
$D_{ab}^{(2)}(w)$ $=\frac{1}{4}r_{a}(m,V)\nu_{b}(\Delta_{c}^{2}\tau_{c}+r_{ac}^{2}w^{2}\tau_{ac}).$ (31)
Lastly, we may close the self-consistency by identifying $m_{ab}$ and $V_{ab}$ with the ensemble average of synapses with drift- and diffusion coefficients Eq. (31). Employing the method from Sec. III.3 in Sec. III, this leads to
$\dot{m}_{ab}$ $=D_{ab}^{(1)}(m_{ab},m,V)$
$\dot{V}_{ab}$ $=2D_{ab}^{(1)\prime}(m_{ab},m,V)V_{ab}$
$\;+\frac{r_{a}(m,V)\nu_{b}}{2}\left[\Delta_{c}^{2}\tau_{c}+r_{ac}^{2}\tau_{ac}\left(V_{ab}+m_{ab}^{2}\right)\right].$ (32)
To summarize, the dynamics of the means and variances of the four populations in training session $k$ is given in terms of eight coupled differential equations Eq. (32) which we integrate numerically. In Fig. 3, we show $m_{ab}$ and $\sqrt{V_{ab}}$ as functions of time. The mean weight of all four populations grows, reflecting that their reversal points are above $m_{0}$.
Figure 3: Population dynamics during a training session of length $T=50$. Mean $m_{ab}$ and standard deviation $\sqrt{V_{ab}}$ of the four populations from Eq. (32) (black line and gray shading) and from simulations (red errorbars). The separate rightmost errorbars schematically illustrate the effect of homeostasis Eq. (34) (we do not consider an actual time for homeostasis). Parameters are neuron: $\mu_{E}=0$, $D_{E}=0.1$, $\mu_{I}=0.5$, $D_{I}=0.05$ network: $N_{E}=4000$, $N_{I}=1000$, $C_{E}=200$, $C_{I}=50$, $(J,g,h)=(0.01,5,2)$ input: $(f_{c},f_{s})=(0.05,0.1)$, $M=200$, $m_{0}=0.05$, $\nu_{\text{hi}}=1$, $\nu_{\text{lo}}=0.1$, $\nu_{s}=64$, $J_{s}=1/80$, STDP: as in Fig. 1.
IV.2 Stationarity on the macroscopic time scale
After evolving with STDP, the synapses undergo homeostasis: the weights are scaled so as to fix the mean weight in $W$ per row to $m_{0}$. This requires rescaling of $W_{ij}\in P_{ab}$ with
$\gamma_{a}=\frac{m_{0}}{f_{c}m_{a1}(T)+(1-f_{c})m_{a0}(T)}.$ (33)
Thus, after the $k$th training session and the subsequent homeostatic process, the population means and variances of $W_{k}$ are
$\gamma_{a}m_{ab}(T),\,\,\text{and}\,\,\gamma_{a}^{2}V_{ab}(T),$ (34)
respectively, these are the separate errorbars in Fig. 3.
While the total average of $W_{k}$,
$N_{E}^{-1}M^{-1}\sum_{ij}W_{k,ij}=m_{0},$ (35)
is fixed by construction, the same is not true for the total variance
$V_{k}=N_{E}^{-1}M^{-1}\sum_{ij}(W_{k,ij}-m_{0})^{2}.$ (36)
However, $V_{k}$ approaches a stationary value due to the interplay of STDP and homeostasis: Assuming the total variance in the beginning of the $k$th training session was $V_{0}$, this fixes the initial condition of all four variances in Eq. (32) to $V_{0}$. The four means and variances after STDP for time $T$ and subsequent homeostasis, Eq. (34), thus parametrically depend on $V_{0}$. Therefore, the total variance of weights in $W$ after the $k$th STDP session and homeostasis, $V_{k}$, depends on $V_{0}$ as well. It can be expressed by the results of integrating Eq. (32) as
$V_{k}(V_{0})$ $=\sum_{ab}^{\ast}[\gamma_{a}^{2}V_{ab}(T)+(\gamma_{a}m_{ab}(T)-m_{0})^{2}],$ (37)
where we defined the weighted population sum
$\sum_{ab}^{\ast}X_{ab}=\sum_{a,b=0}^{1}f_{s}(f_{s}^{-1}-1)^{\delta_{a0}}f_{c}(f_{c}^{-1}-1)^{\delta_{b0}}X_{ab}$ (38)
The asymptotic variance must be stationary
$V_{0}=V_{k}(V_{0}).$ (39)
We solve Eq. (39) with a bisection algorithm. As tested with simulations [see Fig. 4(b)], the variance indeed approaches the result of Eq. (39) to satisfying accuracy.
IV.3 Forgetting on the macroscopic time scale
Figure 4: Dynamics of memory traces and memory capacity. (a) Mean synaptic strength from cue to (non) target neurons after $k$ subsequently stored associations. Simulation results denoted by circles, green for cue-to-target and red for cue-to-non-target, dark colors are averages over 100100 realizations, light colors are single realizations. Theory Eq. (45) (black lines). (b) Variance of synaptic strength from cue to (non) target. Colors as in (a), theory (black lines) from Eq. (46) and the horizontal line is the stationary variance Eq. (39). (c) Fraction of false positives on average over 100 realizations (dark purple circles), for a single realization (light circles) and theory Eq. (53) based on the full drift (black solid line) and neglecting cross-correlations (dotted line). (d,e) Memory capacity Eq. (48) as a function of input sparseness $f_{c}$ (d) and the input width $M$ (e) with (solid) and without (dotted) cross-correlations. Parameters as in Fig. 3, but $\nu_{\text{lo}}=0$.
Here, we take the final steps from neuron- and synapse models to a key property of memory–the memory capacity. To this end, we compute the degradation of a memory trace from which we estimate the fraction of false positives in an attempted recall.
Trace degradation
We assume that the system is the macroscopically stationary state identified above. After the entraining of a specific pattern pair, say $(\boldsymbol{q}_{0},\boldsymbol{p}_{0})$, the weights from $\boldsymbol{q}_{0}$ to $\boldsymbol{p}_{0}$ have mean $K_{11}^{(0)}=\gamma_{1}m_{11}(T)$ and variance $G_{11}^{(0)}=\gamma_{1}^{2}V_{11}(T)$. Analogously, the four distinct populations have mean and variance
$K^{(0)}$ $=\begin{pmatrix}\gamma_{0}m_{00}(T)&\gamma_{0}m_{01}(T)\\ \gamma_{1}m_{10}(T)&\gamma_{1}m_{11}(T)\end{pmatrix}$ (40)
$\text{and}\,\,G^{(0)}$ $=\begin{pmatrix}\gamma_{0}^{2}V_{00}(T)&\gamma_{0}^{2}V_{01}(T)\\ \gamma_{1}^{2}V_{10}(T)&\gamma_{1}^{2}V_{11}(T)\end{pmatrix}$. (41)
To study the course of forgetting, we now derive the evolution of $K^{(k)}$ and $G^{(k)}$ – mean and variance of the populations defined by $(\boldsymbol{q}_{0},\boldsymbol{p}_{0})$ – due to the entrainment of $k$ subsequent pattern pairs. Upon strong self-averaging assumptions, we may assume that each of the four populations is split exactly into four subpopulations, corresponding to the $k$th pattern pair. Further assuming that the shift $m_{ab}(T)-m_{0}$ does not change too much for the different initial $m_{ab}(0)=K_{ab}^{(k-1)}$, we can express the effect of one training session on $K^{(k)}$ as
$K_{ab}^{(k)}$ $=\sum_{a^{\prime}b^{\prime}}^{\ast}\gamma_{a^{\prime}}\left(K_{ab}^{(k-1)}+[m_{a^{\prime}b^{\prime}}(T)-m_{0}]\right).$ (42)
Defining
$\phi$ $=1-\sum_{a=0}^{1}[\delta_{a0}(1-f_{s})+\delta_{a1}f_{s}]\gamma_{a},$
$c$ $=\sum_{ab}^{\ast}\gamma_{a}[m_{ab}(T)-m_{0}],$ (43)
we may express Eq. (42) in the simple format
$K_{ab}^{(k)}=(1-\phi)K_{ab}^{(k-1)}+c.$ (44)
This recursive equation has the solution
$K_{ab}^{(k)}=(1-\phi)^{k}K_{ab}^{(0)}+\phi^{-1}c[1-(1-\phi)^{k}]$ (45)
as one can check by insertion. Thus, for small $\phi$, the elevated weights relax with rate $\phi$; this relaxation rate can be fully expressed through the rescaling factors $\gamma_{a}$. If $\gamma_{a}=1$ (i.e., no rescaling necessary to maintain $m_{0}$), the trace would not degrade at all. For $\gamma_{a}<1$ (compensating for STDP that on average potentiates), it is rather homeostatic downscaling and less overwriting by new memories that limits the lifetime of memory. Asymptotically, all $K_{ab}^{(k)}$ relax to the imposed population mean $m_{0}=\phi^{-1}c$, as one can check by inserting Eq. (33) into Eq. (43).
An analogous argument for the variances leads to
$G_{ab}^{(k)}$ $=\sum_{a^{\prime}b^{\prime}}^{\ast}\gamma_{a^{\prime}}^{2}\Big{[}G_{ab}^{(k-1)}+V_{a^{\prime}b^{\prime}}(T)-V_{0}$
$\;+\left(K_{ab}^{(k-1)}+m_{a^{\prime}b^{\prime}}(T)-m_{0}\right)^{2}\Big{]}-K_{ab}^{(k)2}$ (46)
for which we do not know an explicit solution. Still, Eq. (46) can be evaluated numerically. $K^{(k)}$ and $G^{(k)}$ are shown and compared to simulations in Fig. 4(a,b); within a few hundert patterns, $K^{(k)}$ and $G^{(k)}$ decay to their equilibrium values $m_{0}$ and $V_{0}$, respectively.
Recall
Lastly, we investigate how well target 0 can be recalled when cue 0 is presented after $k$ subsequent training sessions. For simplicity, we only consider the case $\nu_{\text{lo}}=0$, i.e., where non-cue Poisson processes are silent. We follow the approach of [52] and evaluate the dendritic sums: For each post-synaptic neuron, we compute its summed synaptic weight stemming from cue 0 neurons
$s_{i}=\sum_{j}W_{ij}q_{0,j}.$ (47)
The summed weight reflects how strongly each neuron is driven by the cue 0. If the $f_{s}N_{E}$ most strongly driven neurons are the target pattern $\boldsymbol{p}_{0}$, the association is perfectly recalled. We assume that a certain amount of errors can be corrected by a downstream mechanism, e.g., by an attractor network or a perceptron. Therefore, we next derive the fraction $a_{k}^{\text{fp}}$ of false positives in an attempted recall. From this fraction, we define the memory capacity as
$c=\text{argmin}_{k}(|a_{k}^{\text{fp}}-0.5|),$ (48)
which is the number of patterns $c$ for which $a_{c}^{\text{fp}}\leq 0.5$.
Assuming weak correlations between elements of $W$ we can use the central limit theorem to study Eq. (47); the summed weight to target neurons ($a=1$) and non-target neurons ($a=0$) is respectively a Gaussian random variable
$p(s_{i})=\mathcal{N}(s_{i}|f_{c}M\,K_{a1}^{(k)},f_{c}M\,G_{a1}^{(k)}).$ (49)
Marginalizing over all post-synaptic neurons, the distribution of summed weight $s$ is thus
$p(s)=$ $f_{s}\mathcal{N}(s|f_{c}M\,K_{11}^{(k)},f_{c}M\,G_{11}^{(k)})$
$+(1-f_{s})\mathcal{N}(s|f_{c}M\,K_{01}^{(k)},f_{c}M\,G_{01}^{(k)}).$ (50)
The $f_{s}N_{E}$ most active neurons have $s\geq s_{\ast}$, where $s_{\ast}$ is the $1-f_{s}$th quantile of $p$, i.e.,
$\int_{s_{\ast}}^{\infty}ds\,p(s)=f_{s}.$ (51)
This integral can be evaluated to
$\frac{f_{s}}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,M_{11}^{k}}{\sqrt{2f_{c}MV_{11}^{k}}}\right)+\frac{1-f_{s}}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,M_{01}^{k}}{\sqrt{2f_{c}MV_{01}^{k}}}\right)=f_{s},$ (52)
where erfc is the complementary error function. We solve Eq. (52) numerically. Finally, the fraction of false positives is the fraction of probability mass of Eq. (50) above $s_{\ast}$ that is due to non-target neurons
$a_{k}^{\text{fp}}$ $=f_{s}^{-1}\int_{s_{\ast}}^{\infty}ds\,(1-f_{s})\mathcal{N}(s|f_{c}MM_{01}^{k},f_{c}MV_{01}^{k}).$
$=f_{s}^{-1}(1-f_{s})\frac{1}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,M_{01}^{k}}{\sqrt{2f_{c}MV_{01}^{k}}}\right).$ (53)
Eq. (53) is shown in Fig. 4(c). From Eq. (53) one can compute the memory capacity. As well known from Hopfield-like networks, training with sparse patterns is less detrimental to previous memory [53]; this observation remains valid for the spike-coding setup, as shown in Fig. 4(d). Furthermore, we observe in Fig. 4(c-e) that rate-based approximations systematically and substantially overestimate the memory capacity. Our theory that incorporates spike-time–resolving cross-correlations correctly, predicts memory capacities much more faithfully. Lastly, memory capacity grows linearly with the width $M$ of the input layer [Fig. 4(e)].
V Discussion
In this paper we analytically described the stochastic process of the weight of synapses endowed with STDP and driven by stochastic spike trains of pre- and postsynaptic neurons. Specifically, we derived a concise description in form of a Langevin equation that captures the first two Kramers-Moyal coefficients of the true jump process. From this description we computed the dynamics of the mean and variance of an ensemble of synaptic weights. We next studied a training setup where hetero-associations are stored into a feed-forward matrix of synapses endowed with STDP. Through a mean-field theory, we mapped this setup to the single-synapse case above, which led to a quantitative description of the memory lifetime.
We included the effect of pre-post cross-correlations on the synaptic dynamics through an exact link between the cross-correlation and the neuron’s response functions [32]. We found that both for the single-synapse process and for the network process, cross-correlations have a significant impact if synaptic weights are strong enough. The discrepancy between the full solution and a rate-based approximation is particularly prominent at sparse input patterns. Thus, especially in sparse-coding situations, the theory developed here is an important advancement over rate-based approaches.
Two research perspectives on memory and learning have been united in this paper. On the one hand, in the context of STDP, research often focuses on the transient (i.e., non-stationary) dynamics of ensembles of synaptic weights [11, 12, 13, 14, 15, 16, 17, 18, 19]. These dynamics can, e.g., be expressed by continuous-time differential equations of moments of the synaptic weights. The process described in our theory corresponds to this type of dynamics during each training session, which we referred to above as the microscopic time scale. On the other hand, memory is often described as a discrete update process of a weight matrix happening at a macroscopic timescale [54, 55, 56, 51, 57, 58, 52]. Thus, with each discrete time step, a new association is stored such that it can partially overwrite and interfere with previous associations (termed palimpsest [55]). If such a learning scheme incorporates homeostasis, this discrete-time process approaches stationary dynamics. In our process, this second perspective is covered, too: when subsampling the process once after each training session including homeostasis, the weight-matrices $W_{k}$ approach a stationary Markov chain.
Interestingly, we found that the degradation of the memory cue must first be attributed to homeostasis [see Eq. (43) and Eq. (45)] and thus only indirectly occurs due to the storage of new memory. Effectively, the memory degradation with homeostatic plasticity is thus relaxational as opposed to the palimpsest-like forgetting due to overwriting [55], since our representations are large enough for overwriting effects to average out for a while. It would be interesting to clarify experimentally under which circumstances forgetting is rather relaxational or palimpsest-like.
The hetero-associative setting with training induced by exposure, Fig. 1(c), can model memory dynamics on multiple stages. Assuming that the cue and target patterns are sensory representations, our setting models initial retrieval of associations. Additionally, the cue and/ or target patterns could be set by a different synaptic pathway, a teacher. In the context of memory consolidation, the teacher may be a synaptic pathway via the hippocampus which consolidates memory by transferring the association to more stable pathways, as discussed in [59].
The success of the approach at hand opens a vast set of intriguing directions: Previous theory on recurrent plasticity (e.g., Refs. [17, 34]) should be revisited from the present point of view to justify approximations and find corrections. In such recurrent settings, one might need to go beyond the Poissonian-input approximation by applying the results on colored shot noise in [32] and on self-consistent power spectra in [60]. When adding recurrent plasticity to the post-synaptic population in our setup, attracting rate states can arise which certainly impact memory, especially if one considers correlated training patterns.
It has been experimentally reported that synaptic weights fluctuate even in the absence of neural activity [7, 10]; consequently, one could also study an extension of our model in which the intrinsic synaptic noise is taken into account.
Another interesting direction concerns the recent finding that different transfer functions in multi-layer perceptrons give rise to qualitatively different representation (or coding) schemes in the feature layer [61]; correspondingly, it would be quite interesting to study multi-stage feed-forward and locally recurrent networks of spiking neurons under STDP. When exposing the input and readout layer to data, as in the present paper, representations in the hidden layer will arise. It would be interesting to study the statistics of these representations.
Lastly, while for rate-based neural networks the joint neural and synaptic dynamics have been comprehensively described [62], a corresponding theory for spiking (integrate-and-fire) neurons is still missing; the approach presented here is a step in that direction.
Acknowledgements.
We acknowledge helpful discussions with Holger Kantz and Moritz Helias. This work has been funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), SFB1315, project-ID 327654276 to BL and RK.
Appendix A Numerical methods
Simulations of synapses and neurons are implemented in cython [63]. For the LIF neurons, Eq. (4), membrane voltages are integrated with the Euler-Maruyama scheme. In the network, Eq. (5) and Eq. (6), spike detection in one time step results in voltage reset and spike delivery in the same time step. The synaptic model is simulated by integrating Eq. (3) with the Euler scheme. Additionally, whenever the weight is set to a negative value, it is clipped to zero. For the experimentally inspired parameters used in the present paper, this almost never happens. For both Eq. (4) and Eq. (3), the Euler time step is $\Delta t=10^{-4}\tau_{m}$.
Eq. (24) and Eq. (32) are integrated numerically using a Runge-Kutta scheme of order 5(4).
Appendix B Response functions of the LIF neuron
The susceptibilities required in Eq. (17) are the susceptibility of the LIF neuron Eq. (4) to mean- and noise-intensity-modulations, respectively. These have been derived using Fokker-Plank theory in [27],
$\alpha(\Omega)$ $=\frac{ri\Omega/\sqrt{D_{\text{DA}}}}{i\Omega-1}\frac{\mathcal{D}_{i\Omega-1}\left(\frac{\mu_{\text{DA}}-v_{\text{t}}}{\sqrt{D_{\text{DA}}}}\right)-e^{\Delta}\mathcal{D}_{i\Omega-1}\left(\frac{\mu_{\text{DA}}-v_{\text{r}}}{\sqrt{D_{\text{DA}}}}\right)}{\mathcal{D}_{i\Omega}\left(\frac{\mu_{\text{DA}}-v_{\text{t}}}{\sqrt{D_{\text{DA}}}}\right)-e^{\Delta}\mathcal{D}_{i\Omega}\left(\frac{\mu_{\text{DA}}-v_{\text{r}}}{\sqrt{D_{\text{DA}}}}\right)},$ (54)
$\beta(\Omega)$ $=\frac{ri\Omega(i\Omega-1)}{D(2-i\Omega)}\frac{\mathcal{D}_{i\Omega-2}\left(\frac{\mu_{\text{DA}}-v_{\text{t}}}{\sqrt{D_{\text{DA}}}}\right)-e^{\Delta}\mathcal{D}_{i\Omega-2}\left(\frac{\mu_{\text{DA}}-v_{\text{r}}}{\sqrt{D_{\text{DA}}}}\right)}{\mathcal{D}_{i\Omega}\left(\frac{\mu_{\text{DA}}-v_{\text{t}}}{\sqrt{D_{\text{DA}}}}\right)-e^{\Delta}\mathcal{D}_{i\Omega}\left(\frac{\mu_{\text{DA}}-v_{\text{r}}}{\sqrt{D_{\text{DA}}}}\right)},$ (55)
where $\mathcal{D}_{z}(x)$ is Whittaker’s parabolic cylinder function. For the drift coefficient we need to evaluate Eq. (54) and Eq. (55) at imaginary frequencies. This is convenient, as it only requires to evaluate $\mathcal{D}_{z}(x)$ at $z\in\mathbb{R}$. For the figures, we use the implementation $\mathtt{scipy.special.pbdv(z,x)}$.
Appendix C Approximation of $D^{(2)}$
The diffusion coefficient is the second Kramers-Moyal coefficient
$D^{(2)}(w)=\frac{1}{2}\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\left\langle\left[w_{\text{traj}}(t+\Delta t)-w_{\text{traj}}(t)\right]^{2}\right\rangle_{w_{\text{traj}}(t)=w},$ (56)
where $w_{\text{traj}}$ are realizations of Eq. (1). We may express the increment as
$w_{\text{traj}}(t+\Delta t)-w_{\text{traj}}(t)=\int_{t}^{t+\Delta t}dt^{\prime}\dot{w}(t^{\prime})dt^{\prime}.$ (57)
Thus,
$D^{(2)}(w)=\frac{1}{2}\lim_{\tau\rightarrow 0}\frac{1}{\Delta t}\int_{t}^{t+\Delta t}dt^{\prime}\int_{t}^{t+\Delta t}dt^{\prime\prime}\left\langle\dot{w}(t^{\prime})\dot{w}(t^{\prime\prime})\right\rangle_{w_{\text{traj}}(t)=w}.$ (58)
The integration domain is an area of size $\Delta t^{2}$ and we have a prefactor $1/\Delta t$; thus the only terms in the integrand that contribute after taking the limit $\lim_{\Delta t\rightarrow 0}$ are those carrying a $\delta(t^{\prime}-t^{\prime\prime})$. Plugging Eq. (1) into the second moment, we get
$\left\langle\dot{w}(t^{\prime})\dot{w}(t^{\prime\prime})\right\rangle_{w_{\text{traj}}(t)=w}$ $=\Big{\langle}\Big{[}\int_{-\infty}^{0}d\tau\,\kappa(\tau,w)\eta(t^{\prime})x(t^{\prime}+\tau)+\int_{-\infty}^{0}d\tau\,\kappa(-\tau,w)\eta(t^{\prime}+\tau)x(t^{\prime})\Big{]}$
$\times\Big{[}\int_{-\infty}^{0}d\tau\,\kappa(\tau,w)\eta(t^{\prime\prime})x(t^{\prime\prime}+\tau)+\int_{-\infty}^{0}d\tau\,\kappa(-\tau,w)\eta(t^{\prime\prime}+\tau)x(t^{\prime\prime})\Big{]}\Big{\rangle}.$ (59)
Multiplying out the product in Eq. (59) yields four double integrals. Each integral contains a 4-point-correlator $\left\langle\eta(t_{a})\eta(t_{b})x(t_{c})x(t_{d})\right\rangle$, which is in general difficult to evaluate. As an approximation, we assume that we may apply Wick’s theorem for the treatment of this 4-point-correlator even though $(\eta,x)$ is not a Gaussian process. Proceeding with this assumption, we next note that in Eq. (58), only terms carrying a $\delta(t^{\prime}-t^{\prime\prime})$ contribute. Cross-correlations between $x(t_{a})$ and $\eta(t_{b})$ may contain instantaneous parts $\propto\delta(t_{a}-t_{b})$, however, these will eliminate both integrals in Eq. (59) and leave no Dirac delta for the integral in Eq. (58); such contributions thus vanish at $\Delta t\rightarrow 0$. Furthermore, as argued above, constant parts do not contribute. This leaves us with
$\left\langle\eta(t_{a})\eta(t_{b})x(t_{c})x(t_{d})\right\rangle\approx\nu r(w)\delta(t_{a}-t_{b})\delta(t_{c}-t_{d})$ (60)
as the only non-vanishing contribution. The cross products in Eq. (59) produce constant parts and thus do not contribute. Finally, the diagonal products yield
$\left\langle\dot{w}(t^{\prime})\dot{w}(t^{\prime\prime})\Big{|}w_{\text{traj}}(t)=w\right\rangle$ $=\nu r(w)\,\delta(t^{\prime}-t^{\prime\prime})\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)^{2}$ (61)
such that
$D^{(2)}(w)=\frac{1}{2}r(w)\nu\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)^{2}=\frac{1}{4}r(w)\nu(\Delta_{c}^{2}\tau_{c}+r_{ac}^{2}w^{2}\tau_{ac}).$ (62)
Appendix D Mean-field theory of the recurrent network
The drift- and diffusion coefficients of feed-forward synapses in Eq. (31) depend on the input from the recurrent network. To determine its statistics, we follow [42]. The main difference to [42] is that we have three populations, target, non-target, and inhibitory neurons, and that for $h\neq 1$, inhibitory neurons receive stronger excitation than excitatory neurons. The total input to excitatory target and non-target neurons is presented in Eq. (30). The total input to inhibitory neurons is
$\mu_{I}^{\text{tot}}$ $=\mu_{I}+hJC_{E}r_{E}-gJC_{I}r_{I},$
$D_{I}^{\text{tot}}$ $=D_{I}+\frac{1}{2}(hJ)^{2}C_{E}r_{E}+\frac{1}{2}(gJ)^{2}C_{I}r_{I}.$ (63)
Here, the excitatory firing rate is the weighted sum of the target and the non-target neuron’s firing rate
$r_{E}=f_{s}r_{1}+(1-f_{s})r_{0}.$ (64)
Thus, the input to all neurons is determined by the firing rate $r_{1}$, $r_{0}$, and $r_{I}$, which in turn are determined by Eq. (13). This self-consistent set of equations is solved using a damped saddle-point iteration.

### v2

@@
 Abstract
-Learning is based on synaptic plasticity, which drives and is driven by neural activity.
+Learning is based on synaptic plasticity, which affects and is driven by neural activity.
 Because pre- and postsynaptic spiking activity is shaped by randomness, the synaptic weights follow a stochastic process, requiring a probabilistic framework to capture the noisy synaptic dynamics.
@@
 We apply the mean-field theory to characterize forgetting and compute the memory capacity of the setup.
-Figure 1: We investigate memory properties by studying a single synapse dynamics that captures network properties effectively by a mean-field theory.
+Figure 1: We investigate memory properties by studying a single synapse dynamics that effectively captures synaptic population dynamics.
 (a) A Poisson process $\eta$ drives the neuron $x$ through the synapse $w$.
@@
 1(c), which is composed of the synapse- and neuron models above.
-We are interested how hetero-associations can be stored into this network; the necessary training scheme [see Fig.
+We are interested in how hetero-associations can be stored into this network; the necessary training scheme [see Fig.
 1(d)] is explained last.
@@
 1(b), with $\Delta_{c},r_{ac}>0$ determining the amplitudes and $\tau_{c},\tau_{ac}>0$ being the time scales of potentiation and depression, respectively.
-In addition to Eq.
-(1) and Eq.
-(2), we demand that the considered excitatory synapse maintains a positive weight throughout, $w(t)\geq 0$, a necessary condition for Dale’s law, and implemented in our simulations by a clipping boundary condition.
+In addition to Eqs.
+(1) and (2), we demand that the considered excitatory synapse maintains a positive weight throughout, $w(t)\geq 0$, a necessary condition for Dale’s law, and implemented in our simulations by a clipping boundary condition.
 The plasticity rule Eq.
@@
 $\dot{A}_{\text{post}}$ $=-\tau_{ac}^{-1}A_{\text{post}}+x.$ (3)
-This embedding is useful for simulations as it is makes the system local in time.
+This embedding is useful for simulations as it makes the system local in time.
 To avoid switch-on effects in Eq.
@@
 We here introduce a neuron model that is driven (i) by Gaussian white noise and (ii) by a single presynaptic neuron through a single synapse.
-The dynamics of this synapse will be studied in Sec.
+The dynamics of this single synapse will be studied in Sec.
 III.
@@
 A layer of $M$ Poisson processes $\eta_{i}$ (gray circles) drives a recurrent network of $N$ LIF neurons [42] (red and blue circles) through a plastic weight matrix $W$.
-The recurrent network consists of $N_{E}$ excitatory and $N_{I}$ inhibitory neurons, among which only the excitatory neurons are targeted by $W$.
+The recurrent network consists of $N_{E}$ excitatory (red) and $N_{I}$ inhibitory (blue) neurons, among which only the excitatory neurons are targeted by $W$.
 Addidionally, each neuron has exactly $C_{E}$ incoming excitatory connections from the recurrent network and $C_{I}$ incoming inhibitory connections with static weights $J_{EE}=J$, $J_{EI}=J_{II}=-gJ$, $J_{IE}=hJ$, parameterized by $J>0$, $g>0$, and $h>0$.
-Most input Poisson processes have a low firing rate $\nu_{\text{lo}}$, however, in each training step, a subset of $f_{c}M$ Poisson generators, represented by the binary vector $\boldsymbol{q}\in\left\{0,1\right\}^{M}$ fires with increased rate $\nu_{\text{hi}}$.
+Most input Poisson processes have a low firing rate $\nu_{\text{lo}}$, however, in each training step, a subset of $f_{c}M$ Poisson generators, represented by the ones in binary vector $\boldsymbol{q}\in\left\{0,1\right\}^{M}$ fires with increased rate $\nu_{\text{hi}}$.
 Simultaneously, a subset of $f_{s}N_{E}$ excitatory neurons, represented by the binary vector $\boldsymbol{p}\in\left\{0,1\right\}^{N_{E}}$, receives direct Poissonian input $\gamma_{i}$ with rate $\nu_{s}$ and static weight $J_{s}$; the other $(1-f_{s})N_{E}$ excitatory neurons do not receive additional input.
@@
 Trajectories of $w$ can be obtained in simulations, for instance, for the window function Eq.
-(2), one has to integrate both stochastic equations, Eq.
-(3) and Eq.
-(4).
+(2), one has to integrate both stochastic equations, Eqs.
+(3) and (4).
 The synaptic weight performs a jump process Fig.
@@
 Within a diffusion approximation of the input to the LIF neuron
-$\mu+\sqrt{2D}\xi(t)+w\eta(t)$
-$\approx$ $\mu_{\text{DA}}(\nu)+\sqrt{2D_{\text{DA}}(\nu)}\xi(t),$ (12)
-where $\mu_{\text{DA}}(\nu)=\mu+w\nu$ and $D_{\text{DA}}(\nu)=D+w^{2}\nu/2$, the instantaneous firing rate is given by [40]
-$r^{-1}=\int_{\frac{v_{r}-\mu_{\text{DA}}}{\sqrt{2D_{\text{DA}}}}}^{\frac{v_{t}-\mu_{\text{DA}}}{\sqrt{2D_{\text{DA}}}}}ds\,e^{s^{2}}[1+\text{erf}(s)],$ (13)
+$\mu+\sqrt{2D}\xi(t)+w\eta(t)\approx\mu_{\text{DA}}+\sqrt{2D_{\text{DA}}}\xi(t),$ (12)
+where $\mu_{\text{DA}}=\mu+w\nu$ and $D_{\text{DA}}=D+w^{2}\nu/2$, the instantaneous firing rate is given by [40]
+$\frac{1}{r(w)}=\int_{\frac{v_{r}-\mu_{\text{DA}}}{\sqrt{2D_{\text{DA}}}}}^{\frac{v_{t}-\mu_{\text{DA}}}{\sqrt{2D_{\text{DA}}}}}ds\,e^{s^{2}}[1+\text{erf}(s)],$ (13)
 see [48] for an efficient and stable evaluation of Eq. (13). For the STDP kernel in Eq. (2), the first, firing-rate–based, part in Eq. (11) is
@@
 For non-infinitesimal $\Delta t$ long enough (specifically, in Fig.
-2(c), $\Delta t=10$), the diffusion is captured by the Langevin dynamics based on Eq.
-(17) and Eq.
-(19), as we derive in the next section.
+2(c), $\Delta t=10$), the diffusion is captured by the Langevin dynamics based on Eqs.
+(17) and (19), as we derive in the next section.
 III.3 Mid- and long-term evolution of an ensemble of synapses
@@
 Here, we study the mean and variance of an ensemble of synapses on finite times.
-We provide a technically detailed analysis because the results of this section are the main building blocks of the mean-field theory of learning discussed in Sec.
-IV.
+We provide a technically detailed analysis because the results of this section are the main building blocks of the mean-field theory of learning discussed in the next section (Sec.
+IV).
 Although the dynamics can admit a stationary solution [see e.g.
@@
 $A(t)$ $=\int_{0}^{t}2D^{(1)\prime}[m(t^{\prime})]dt^{\prime}+\frac{t}{2}r\nu r_{ac}^{2}\tau_{ac}.$ (26)
-Eq.
-(24) and Eq.
-(26) can be efficiently evaluated to obtain the mean and variance of an ensemble of synapses.
+Equations (24) and (26) can be efficiently evaluated to obtain the mean and variance of an ensemble of synapses.
 These equations are thus useful formulations of the microscopic (i.e., single-synapse) model, which we leverage in the next section to derive macroscopic (i.e., population) dynamics.
-We also use Eq.
-(24) and Eq.
-(26) to verify the second moment of the stochastic increment over a finite time window Fig.
+We also use Eqs.
+(24) and (26) to verify the second moment of the stochastic increment over a finite time window Fig.
 2(c); i.e., due to the invalidity of the infinitesimal diffusion coefficient, we predict the finite-time diffusion coefficient as put forward in [50].
@@
 In this section we study the network and learning scheme sketched in Fig.
-1(c,d) and introduced in Sec.
-II.3 and Sec.
-II.4 in Sec.
-II.
+1(c,d) (introduced in Secs.
+II.3 and II.4).
 The dynamics can now be viewed on two time scales: First, on the discrete macroscopic time scale (pattern index $k$, going from one training session to the next), the data that the network is exposed to is exchanged.
@@
 (30), the postsynaptic neurons are, within the employed approximation, effectively decoupled apart from their common dependence on the mean fields $m_{ab}$, $V_{ab}$, $r_{E}$, and $r_{I}$, which we now determine self-consistently.
-Thus, assuming knowledge of the numerical values of the mean fields, the single-neuron statistics is readily determined and we can proceed as in Sec.
+Thus, assuming knowledge of the numerical values of the mean fields, the single-neuron statistics is readily determined and we can proceed as for the single synapse in Sec.
 III: The drift- and diffusion coefficients of the four distinct synaptic populations are, respectively
@@
 Employing the method from Sec.
-III.3 in Sec.
-III, this leads to
+III.3, this leads to
 $\dot{m}_{ab}$ $=D_{ab}^{(1)}(m_{ab},m,V)$
@@
 Figure 4: Dynamics of memory traces and memory capacity.
-(a) Mean synaptic strength from cue to (non) target neurons after $k$ subsequently stored associations.
+(a) Mean synaptic strength from cue to (non-)target neurons after $k$ subsequently stored associations.
 Simulation results denoted by circles, green for cue-to-target and red for cue-to-non-target, dark colors are averages over 100100 realizations, light colors are single realizations.
@@
 (45) (black lines).
-(b) Variance of synaptic strength from cue to (non) target.
+(b) Variance of synaptic strength from cue to (non-)target.
 Colors as in (a), theory (black lines) from Eq.
@@
 To this end, we compute the degradation of a memory trace from which we estimate the fraction of false positives in an attempted recall.
-Trace degradation
-We assume that the system is the macroscopically stationary state identified above.
+IV.3.1 Trace degradation
+We assume that the system is in the macroscopically stationary state identified above.
 After the entraining of a specific pattern pair, say $(\boldsymbol{q}_{0},\boldsymbol{p}_{0})$, the weights from $\boldsymbol{q}_{0}$ to $\boldsymbol{p}_{0}$ have mean $K_{11}^{(0)}=\gamma_{1}m_{11}(T)$ and variance $G_{11}^{(0)}=\gamma_{1}^{2}V_{11}(T)$.
@@
 $K^{(k)}$ and $G^{(k)}$ are shown and compared to simulations in Fig.
-4(a,b); within a few hundert patterns, $K^{(k)}$ and $G^{(k)}$ decay to their equilibrium values $m_{0}$ and $V_{0}$, respectively.
-Recall
+4(a,b); within a few hundred patterns, $K^{(k)}$ and $G^{(k)}$ decay to their equilibrium values $m_{0}$ and $V_{0}$, respectively.
+IV.3.2 Recall
 Lastly, we investigate how well target 0 can be recalled when cue 0 is presented after $k$ subsequent training sessions.
@@
 This integral can be evaluated to
-$\frac{f_{s}}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,M_{11}^{k}}{\sqrt{2f_{c}MV_{11}^{k}}}\right)+\frac{1-f_{s}}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,M_{01}^{k}}{\sqrt{2f_{c}MV_{01}^{k}}}\right)=f_{s},$ (52)
+$\frac{f_{s}}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,K_{11}^{(k)}}{\sqrt{2f_{c}MG_{11}^{(k)}}}\right)+\frac{1-f_{s}}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,K_{01}^{(k)}}{\sqrt{2f_{c}MG_{01}^{(k)}}}\right)=f_{s},$ (52)
 where erfc is the complementary error function.
@@
 (50) above $s_{\ast}$ that is due to non-target neurons
-$a_{k}^{\text{fp}}$ $=f_{s}^{-1}\int_{s_{\ast}}^{\infty}ds\,(1-f_{s})\mathcal{N}(s|f_{c}MM_{01}^{k},f_{c}MV_{01}^{k}).$
-$=f_{s}^{-1}(1-f_{s})\frac{1}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,M_{01}^{k}}{\sqrt{2f_{c}MV_{01}^{k}}}\right).$ (53)
+$a_{k}^{\text{fp}}$ $=f_{s}^{-1}\int_{s_{\ast}}^{\infty}ds\,(1-f_{s})\mathcal{N}(s|f_{c}MK_{01}^{(k)},f_{c}MG_{01}^{(k)})$
+$=f_{s}^{-1}(1-f_{s})\frac{1}{2}\text{erfc}\left(\frac{s_{\ast}-f_{c}M\,K_{01}^{(k)}}{\sqrt{2f_{c}MG_{01}^{(k)}}}\right).$ (53)
 Eq.
@@
 (4), membrane voltages are integrated with the Euler-Maruyama scheme.
-In the network, Eq.
-(5) and Eq.
-(6), spike detection in one time step results in voltage reset and spike delivery in the same time step.
+In the network, Eqs.
+(5) and (6), spike detection in one time step results in voltage reset and spike delivery in the same time step.
 The synaptic model is simulated by integrating Eq.
@@
 (3), the Euler time step is $\Delta t=10^{-4}\tau_{m}$.
-Eq. (24) and Eq. (32) are integrated numerically using a Runge-Kutta scheme of order 5(4).
+Equations (24) and (32) are integrated numerically using a Runge-Kutta scheme of order 5(4).
 Appendix B Response functions of the LIF neuron
@@
 (4) to mean- and noise-intensity-modulations, respectively.
-These have been derived using Fokker-Plank theory in [27],
+These have been derived using Fokker-Planck theory in [27],
 $\alpha(\Omega)$ $=\frac{ri\Omega/\sqrt{D_{\text{DA}}}}{i\Omega-1}\frac{\mathcal{D}_{i\Omega-1}\left(\frac{\mu_{\text{DA}}-v_{\text{t}}}{\sqrt{D_{\text{DA}}}}\right)-e^{\Delta}\mathcal{D}_{i\Omega-1}\left(\frac{\mu_{\text{DA}}-v_{\text{r}}}{\sqrt{D_{\text{DA}}}}\right)}{\mathcal{D}_{i\Omega}\left(\frac{\mu_{\text{DA}}-v_{\text{t}}}{\sqrt{D_{\text{DA}}}}\right)-e^{\Delta}\mathcal{D}_{i\Omega}\left(\frac{\mu_{\text{DA}}-v_{\text{r}}}{\sqrt{D_{\text{DA}}}}\right)},$ (54)
@@
 where $\mathcal{D}_{z}(x)$ is Whittaker’s parabolic cylinder function.
-For the drift coefficient we need to evaluate Eq.
-(54) and Eq.
-(55) at imaginary frequencies.
+For the drift coefficient we need to evaluate Eqs.
+(54) and (55) at imaginary frequencies.
 This is convenient, as it only requires to evaluate $\mathcal{D}_{z}(x)$ at $z\in\mathbb{R}$.
@@
 as the only non-vanishing contribution. The cross products in Eq. (59) produce constant parts and thus do not contribute. Finally, the diagonal products yield
-$\left\langle\dot{w}(t^{\prime})\dot{w}(t^{\prime\prime})\Big{|}w_{\text{traj}}(t)=w\right\rangle$ $=\nu r(w)\,\delta(t^{\prime}-t^{\prime\prime})\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)^{2}$ (61)
+$\left\langle\dot{w}(t^{\prime})\dot{w}(t^{\prime\prime})\right\rangle_{w_{\text{traj}}(t)=w}$ $=\nu r(w)\,\delta(t^{\prime}-t^{\prime\prime})\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)^{2}$ (61)
 such that
-$D^{(2)}(w)=\frac{1}{2}r(w)\nu\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)^{2}=\frac{1}{4}r(w)\nu(\Delta_{c}^{2}\tau_{c}+r_{ac}^{2}w^{2}\tau_{ac}).$ (62)
+$D^{(2)}(w)$ $=\frac{1}{2}r(w)\nu\int_{-\infty}^{\infty}d\tau\,\kappa(\tau,w)^{2}$
+$=\frac{1}{4}r(w)\nu(\Delta_{c}^{2}\tau_{c}+r_{ac}^{2}w^{2}\tau_{ac}).$ (62)
 Appendix D Mean-field theory of the recurrent network
@@
 (13).
-This self-consistent set of equations is solved using a damped saddle-point iteration.
+This self-consistent set of equations is solved using a damped saddle-point